OpenStudy (anonymous):

algebra help..... find the standard form for these points (5,3) , (4,5)

OpenStudy (anonymous):

Standard form : \[y = ax + b\]Slope a : \[a = \frac{B_y-A_y}{B_x-A_x}= \frac{3-5}{5-4}=-2\]b:\[-2\times5+b=3\]\[b=13\]Thus\[y=-2x+13\] (Just ask if you don't understand something)

OpenStudy (anonymous):

Standard form should be x's + y's = number so this should be writtne 2x + y = 13